OpenStudy (anonymous):

a 12% simple interest loan was taken out and will be repaid in 18 months , with a payoff of $1062. how much is being borrowed?

OpenStudy (gw2011):

The borrowed amount is $900. Here is how it was calculated: I=(P)(R)(T) where I=Interest, P=Principal, and T=Time $1062-P=(P)(0.12)(1.5) 1062-P=0.18P 1062=0.18P+P 1062=1.18P and P=$900